Stock price of $42, earnings of $2.12 per share during the last twelve months, forecasted earnings of $2.84 over the following year, and average earnings growth forecast of 12.5% per year for the next five years. What is this stock's PEG, rounded to one decimal place?
Answer #1

There are two methods of calculating the PEG ratio i.e. Forward PEG and Trailing PEG

For Trailing PEG we need the past growth rate, as there is no information provided of past growth and no data to calculate it, we cannot calculate Trailing PEG.

We will only Calculate Forward PEG ratio based on the data given to us.

PEG Ratio is calculated as, PE ratio divided by Earning growth rate;

PEG Ratio = P/E Ratio / Earnings Growth Rate

PE ratio = Stock Price divided by Earning per share (EPS),

P/E = Stock Price/EPS

(as we are calculating the Forward PEG we will use forecasted earning which is given as $ 2.84)

P/E = $42/$2.84

P/E = 14.78873

(we are given average earnings growth forecast of 12.5% per year for the next five years)

Substituting it in PEG formula;

Forward PEG = P/E / Earning growth rate

= 14.78873/ 12.5

= 1.1830 (rounded to 1 decimal)

= 1.2

(Shown calculation in Excel with formulas also)

B C D $ $ 42.00 2.84 1 2 Stock Price 3 Forecasted earning 4 5 P/E = 6 7 Earning growth rate 14.78873239 =B2/B3 12.50 8. 1.2 °

So the Forward PEG ratio is 1.2